# Learning Disentangled Identifiers for Action-Customized Text-to-Image Generation

### v3

@@ 4.1 Preliminaries
Our study is based on the Stable Diffusion (SD) [22] model, which is considered to be the public state-of-the-art text-to-image generator. Specifically, to operate the diffusion process [6] in a low-dimensional latent space, SD employs a hierarchical VAE that consists of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder $\mathcal{E}$ is tasked with encoding the given image $\mathbf{x}$ into latent features $\mathbf{z}$, and the decoder $\mathcal{D}$ reconstructs the image $\widehat{\mathbf{x}}$ from the latent, i.e., $\widehat{\mathbf{x}}=\mathcal{D}(\mathbf{z})=\mathcal{D}(\mathcal{E}(\mathbf{x}))$. To control the generation with the textual conditions, given the noisy latent $\mathbf{z}_{t}$, current time step $t$ and text tokens $\mathbf{y}$, a conditional U-Net [23] denoiser is trained to predict the noise $\epsilon$ added to the latent $\mathbf{z}$:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}\sim\mathcal{E}(\mathbf{x}),\mathbf{y},\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{z}_{t},t,\mathbf{y}\right)\right\|_{2}^{2}\right],$ (1)
where $\mathbf{y}$ is obtained by feeding the prompt into a CLIP [17] text encoder. During inference, the pre-trained SD first samples a latent $\mathbf{z}_{T}$ from the standard normal distribution $\mathcal{N}(0,1)$. Iteratively, $\mathbf{z}_{t-1}$ can be obtained by removing noise from $\mathbf{z}_{t}$ conditioned on $\mathbf{y}$. After the final denoising step, the latent $\mathbf{z}_{0}$ is mapped to generate an image $\widehat{\mathbf{x}}$ with the decoder $\mathcal{D}$.
@@ 4.2 Action-Disentangled Identifier (ADI)
Given exemplar images that all contain a specific entity, existing subject-driven inversion methods [5, 9] learn to represent the entity as an identifier token $v\in\mathbb{R}^{d}$. And the learned $v$ can then be employed in text prompts to produce diverse and novel images, where the entity can be generated with different contexts. In this paper, we continue the vein of capturing the common action in exemplar images by finding the optimal identifiers. An overview of our proposed ADI is illustrated in Fig. 3.
Expanding Semantic Inversion. To overcome the preference to low-level appearance features, we apply layer-wise identifier tokens to increase the accommodation of various features. Specifically, for the $l$-th layer where $l\in[1,L]$ and $L$ is the number of cross-attention layers in the T2I model, a new identifier token $v_{l}\in\mathbb{R}^{d}$ is initialized. Feeding the prompt with $v_{l}$ into the text encoder, the output tokens $\mathbf{y}_{l}$ control the update of the latents in the $l$-th layer, thus influencing the generation of the visual content. And the learned tokens from all layers can form a token set $\mathcal{V}$, which can then be paired with different subjects for generation. Rather than having a single identifier token take on the responsibility of reconstruction, having separate identifiers at different layers effectively ensures that more features are converted, including the action-related features we care about.
Learning Gradient Mask with Context-Different Pair. The next step is to prevent the identifiers from inverting features that are not relevant to the action and thus contaminating the subsequent image generation. Given $\mathbf{x}^{(a,c)}\in\mathcal{X}$ as an anchor sample, where $a$ denotes the specific action, and $c$ denotes the action-agnostic context contained in the image including human appearance and background, we can randomly sample another image $\mathbf{x}^{(a,\overline{c})}$ from $\mathcal{X}$, where $\overline{c}$ represents that the context is different from $c$. Taking the context-different pair $\mathbf{x}^{(a,c)}$ and $\mathbf{x}^{(a,\overline{c})}$ as the input, we can calculate two gradients of the denoising loss $\mathcal{L}$ with respect to the identifier token $v$:
$g^{(a,c)}=\frac{\partial\mathcal{L}^{(a,c)}}{\partial v},$ (2)
$g^{(a,\overline{c})}=\frac{\partial\mathcal{L}^{(a,\overline{c})}}{\partial v}.$ (3)
@@
$\bigtriangleup g^{\overline{c}}=|g^{(a,c)}-g^{(a,\overline{c})}|,$ (4)
where the semantic channels with a small difference can be regarded as action-related channels of the action $a$, which are expected to be preserved. Specifically, we sort the difference from the largest to the smallest, taking the value at $\beta$ percent $\gamma^{\beta}$ as a threshold. In other words, $\beta$of the channels are masked. Then, the mask that shares the same dimension as $v$ can be calculated. For the $k$-th channel,
$m^{\overline{c}}_{k}=\left\{\begin{array}[]{lc}0,&\bigtriangleup g^{\overline{c}}_{k}\geqslant\gamma^{\beta}\\ 1,&\bigtriangleup g^{\overline{c}}_{k}<\gamma^{\beta}\\ \end{array}.\right.$ (7)
By overwriting the mask to the gradient of the anchor sample, the action-related knowledge is preserved and incorporated into the update of $v$, while the updates on action-agnostic channels are ignored. Note that since the specific visual invariance about the action changes slightly depending on the sample pair, the masked channels may not be exactly the same each time. Furthermore, both samples use the prompt of the anchor sample $\mathbf{x}^{(a,c)}$ when calculating the gradients. Since the visual context of $\mathbf{x}^{(a,\overline{c})}$ is inconsistent with the description in the prompt, the reconstruction loss favours larger gradients in the context-related channels. In this way, the action-related channels found through the threshold will be more accurate.
@@
Learning Gradient Mask with Action-Different Pair. Although the context-different pairs have the same action semantics, there may be differences in the visualization of the actions, and therefore the channels associated with the most representative action features do not necessarily have a smaller gradient difference. Since learning the gradient mask with the context-different only pair is not stable and effective enough, we also construct action-different pairs to generate the gradient mask from another perspective. For each sample $\mathbf{x}^{(a,c)}$ in $\mathcal{X}$, we can use it to quickly train a subject-driven customization model (e.g., DreamBooth) that effectively inverts the most of the low-level context information. Therefore, by filling the prompt template of $\mathbf{x}^{(a,c)}$ with action descriptions that are different from $a$, the trained customization model can generate various action images as $\mathcal{X}^{(\overline{a},c)}$. Note that this step is not necessary if the users can compile a dataset of varied actions by the same individual using pre-captured images. However, the data collection is usually arduous and lengthy, making fast training of a subject-driven customization a more convenient solution. Due to the one-shot training and the concise text, the generated images $\mathcal{X}^{(\overline{a},c)}$ may not be consistent with the action descriptions, or the context may differ from the original $\mathbf{x}^{(a,c)}$, but in practice, we have found that the quality is sufficient to diversify the action variation. In this way, when $\mathbf{x}^{(a,c)}$ is sampled during training, we can randomly sample a image $\mathbf{x}^{(\overline{a},c)}$ from $\mathcal{X}^{(\overline{a},c)}$ to construct the action-different pair. And the gradient of $\mathbf{x}^{(\overline{a},c)}$ with respect to the token $v$ can be calculated as
$g^{(\overline{a},c)}=\frac{\partial\mathcal{L}^{(\overline{a},c)}}{\partial v}.$ (8)
Similarly, both samples use the prompt of the anchor sample $\mathbf{x}^{(a,c)}$. We can also calculate the absolute value of the gradient difference:
$\bigtriangleup g^{\overline{a}}=|g^{(a,c)}-g^{(\overline{a},c)}|,$ (9)
@@
$m^{\overline{a}}_{k}=\left\{\begin{array}[]{lc}0,&\bigtriangleup g^{\overline{a}}_{k}<\lambda^{\beta}\\ 1,&\bigtriangleup g^{\overline{a}}_{k}\geqslant\lambda^{\beta}\\ \end{array},\right.$ (12)
@@
Merging Gradient Masks for Context. Due to the noise introduced by context variations, identifying action-relevant channels using only context-different or action-different pairs would be difficult and unreliable. As an evidence, the average overlap rate of channels preserved by both masks at each training step is 30.26%. Therefore, given the input triple =$\mathcal{I}=\left\{\mathbf{x}^{(a,\overline{c})},\mathbf{x}^{(a,c)},\mathbf{x}^{(\overline{a},c)}\right\}$, we can merge the two obtained masks $m^{\overline{a}}$ and $m^{\overline{c}}$ to get the final context mask $m$. In practice, we keep only the intersection of the unmasked channels as unmasked, as we find this merging strategy performs better. Formally, we have
$m=m^{\overline{c}}\cap{m}^{\overline{a}},$ (13)
@@
$\widetilde{g}^{(a,c)}=m\odot g^{(a,c)}.$ (14)
Note that the masked gradient $\widetilde{g}^{(a,c)}$, where action-agnostic channels are considered to be masked, is the only gradient used to update $v$. Therefore, our identifiers can adequately invert action-related features.
@@ 5.5 Further Analysis
Impact of Gradient Mask Merging Strategy. As shown in Eq. 13, ADI takes the intersection of the two gradient masks as the default merging strategy. We compare this with selecting the union of the two masks, and illustrate the generation results in Fig. 7. Since only channels that are preserved on both masks are updated, taking the intersection can effectively filter out action-agnostic features, leading to better customization of the actions. In contrast, taking the intersection may dilute the most representative action features due to the preserved context information.

### v5

@@ 4.1 Preliminaries
Our study is based on the Stable Diffusion (SD) [22] model, which is considered to be the public state-of-the-art text-to-image generator. Specifically, to operate the diffusion process [6] in a low-dimensional latent space, SD employs a hierarchical VAE that consists of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder $\mathcal{E}$ is tasked with encoding the given image $x$ into latent features $\mathbf{z}$, and the decoder $\mathcal{D}$ reconstructs the image $\widehat{x}$ from the latent, i.e., $\widehat{x}=\mathcal{D}(\mathbf{z})=\mathcal{D}(\mathcal{E}(x))$. To control the generation with the textual conditions, given the noisy latent $\mathbf{z}_{t}$, current time step $t$ and text tokens $\mathbf{y}$, a conditional U-Net [23] denoiser is trained to predict the noise $\epsilon$ added to the latent $\mathbf{z}$:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}\sim\mathcal{E}(x),\mathbf{y},\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{z}_{t},t,\mathbf{y}\right)\right\|_{2}^{2}\right],$ (1)
where $\mathbf{y}$ is obtained by feeding the prompt into a CLIP [17] text encoder. During inference, the pre-trained SD first samples a latent $\mathbf{z}_{T}$ from the standard normal distribution $\mathcal{N}(0,1)$. Iteratively, $\mathbf{z}_{t-1}$ can be obtained by removing noise from $\mathbf{z}_{t}$ conditioned on $\mathbf{y}$. After the final denoising step, the latent $\mathbf{z}_{0}$ is mapped to generate an image $\widehat{x}$ with the decoder $\mathcal{D}$.
@@ 4.2 Action-Disentangled Identifier (ADI)
Given exemplar images that all contain a specific entity, existing subject-driven inversion methods [5, 9] learn to represent the entity as an identifier token $\mathbf{v}\in\mathbb{R}^{d}$. And the learned $\mathbf{v}$ can then be employed in text prompts to produce diverse and novel images, where the entity can be generated with different contexts. In this paper, we continue the vein of capturing the common action in exemplar images by finding the optimal identifiers. An overview of our proposed ADI is illustrated in Fig. 3.
Expanding Semantic Inversion. To overcome the preference to low-level appearance features, we apply layer-wise identifier tokens to increase the accommodation of various features. Specifically, for the $l$-th layer where $l\in[1,L]$ and $L$ is the number of cross-attention layers in the T2I model, a new identifier token $\mathbf{v}_{l}\in\mathbb{R}^{d}$ is initialized. Feeding the prompt with $\mathbf{v}_{l}$ into the text encoder, the output tokens $\mathbf{y}_{l}$ control the update of the latents in the $l$-th layer, thus influencing the generation of the visual content. And the learned tokens from all layers can form a token set $\mathcal{V}$, which can then be paired with different subjects for generation. Rather than having a single identifier token take on the responsibility of reconstruction, having separate identifiers at different layers effectively ensures that more features are converted, including the action-related features we care about.
Learning Gradient Mask with Context-Different Pair. The next step is to prevent the identifiers from inverting features that are not relevant to the action and thus contaminating the subsequent image generation. Given $x^{(a,c)}\in\mathcal{X}$ as an anchor sample, where $a$ denotes the specific action, and $c$ denotes the action-agnostic context contained in the image including human appearance and background, we can randomly sample another image $x^{(a,\overline{c})}$ from $\mathcal{X}$, where $\overline{c}$ represents that the context is different from $c$. Taking the context-different pair $x^{(a,c)}$ and $x^{(a,\overline{c})}$ as the input, we can calculate two gradients of the denoising loss $\mathcal{L}$ with respect to the identifier token $\mathbf{v}$:
$\mathbf{g}^{(a,c)}=\frac{\partial\mathcal{L}^{(a,c)}}{\partial\mathbf{v}},$ (2)
$\mathbf{g}^{(a,\overline{c})}=\frac{\partial\mathcal{L}^{(a,\overline{c})}}{\partial\mathbf{v}}.$ (3)
@@
$\bigtriangleup\mathbf{g}^{\overline{c}}=|\mathbf{g}^{(a,c)}-\mathbf{g}^{(a,\overline{c})}|,$ (4)
where the semantic channels with a small difference can be regarded as action-related channels of the action $a$, which are expected to be preserved. Specifically, we sort the difference from the largest to the smallest, taking the value at $\beta$ percent $\gamma^{\beta}$ as a threshold. In other words, $\beta$of the channels are masked. Then, the mask that shares the same dimension as $\mathbf{v}$ can be calculated. For the $k$-th channel,
$\mathbf{m}^{\overline{c}}_{k}=\left\{\begin{array}[]{lc}0,&\bigtriangleup\mathbf{g}^{\overline{c}}_{k}\geqslant\gamma^{\beta}\\ 1,&\bigtriangleup\mathbf{g}^{\overline{c}}_{k}<\gamma^{\beta}\\ \end{array}.\right.$ (7)
By overwriting the mask to the gradient of the anchor sample, the action-related knowledge is preserved and incorporated into the update of $\mathbf{v}$, while the updates on action-agnostic channels are ignored. Note that since the specific visual invariance about the action changes slightly depending on the sample pair, the masked channels may not be exactly the same each time. Furthermore, both samples use the prompt of the anchor sample $x^{(a,c)}$ when calculating the gradients. Since the visual context of $x^{(a,\overline{c})}$ is inconsistent with the description in the prompt, the reconstruction loss favours larger gradients in the context-related channels. In this way, the action-related channels found through the threshold will be more accurate.
@@
Learning Gradient Mask with Action-Different Pair. Although the context-different pairs have the same action semantics, there may be differences in the visualization of the actions, and therefore the channels associated with the most representative action features do not necessarily have a smaller gradient difference. Since learning the gradient mask with the context-different only pair is not stable and effective enough, we also construct action-different pairs to generate the gradient mask from another perspective. For each sample $x^{(a,c)}$ in $\mathcal{X}$, we can use it to quickly train a subject-driven customization model (e.g., DreamBooth) that effectively inverts the most of the low-level context information. Therefore, by filling the prompt template of $x^{(a,c)}$ with action descriptions that are different from $a$, the trained customization model can generate various action images as $\mathcal{X}^{(\overline{a},c)}$. Note that this step is not necessary if the users can compile a dataset of varied actions by the same individual using pre-captured images. However, the data collection is usually arduous and lengthy, making fast training of a subject-driven customization a more convenient solution. Due to the one-shot training and the concise text, the generated images $\mathcal{X}^{(\overline{a},c)}$ may not be consistent with the action descriptions, or the context may differ from the original $x^{(a,c)}$, but in practice, we have found that the quality is sufficient to diversify the action variation. In this way, when $x^{(a,c)}$ is sampled during training, we can randomly sample a image $x^{(\overline{a},c)}$ from $\mathcal{X}^{(\overline{a},c)}$ to construct the action-different pair. And the gradient of $x^{(\overline{a},c)}$ with respect to the token $\mathbf{v}$ can be calculated as
$\mathbf{g}^{(\overline{a},c)}=\frac{\partial\mathcal{L}^{(\overline{a},c)}}{\partial\mathbf{v}}.$ (8)
Similarly, both samples use the prompt of the anchor sample $x^{(a,c)}$. We can also calculate the absolute value of the gradient difference:
$\bigtriangleup\mathbf{g}^{\overline{a}}=|\mathbf{g}^{(a,c)}-\mathbf{g}^{(\overline{a},c)}|,$ (9)
@@
$\mathbf{m}^{\overline{a}}_{k}=\left\{\begin{array}[]{lc}0,&\bigtriangleup\mathbf{g}^{\overline{a}}_{k}<\lambda^{\beta}\\ 1,&\bigtriangleup\mathbf{g}^{\overline{a}}_{k}\geqslant\lambda^{\beta}\\ \end{array},\right.$ (12)
@@
Merging Gradient Masks for Context. Due to the noise introduced by context variations, identifying action-relevant channels using only context-different or action-different pairs would be difficult and unreliable. As an evidence, the average overlap rate of channels preserved by both masks at each training step is 30.26%. Therefore, given the input triple =$\mathcal{I}=\left\{x^{(a,\overline{c})},x^{(a,c)},x^{(\overline{a},c)}\right\}$, we can merge the two obtained masks $\mathbf{m}^{\overline{a}}$ and $\mathbf{m}^{\overline{c}}$ to get the final context mask $\mathbf{m}$. In practice, we keep only the intersection of the unmasked channels as unmasked, as we find this merging strategy performs better. Formally, we have
$\mathbf{m}=\mathbf{m}^{\overline{c}}\cap{\mathbf{m}}^{\overline{a}},$ (13)
@@
$\widetilde{\mathbf{g}}^{(a,c)}=\mathbf{m}\odot\mathbf{g}^{(a,c)}.$ (14)
Note that the masked gradient $\widetilde{\mathbf{g}}^{(a,c)}$, where action-agnostic channels are considered to be masked, is the only gradient used to update $\mathbf{v}$. Therefore, our identifiers can adequately invert action-related features.
@@ 5.5 Further Analysis
Impact of Gradient Mask Merging Strategy. As shown in Eq. 13, ADI takes the intersection of the two gradient masks as the default merging strategy. We compare this with selecting the union of the two masks, and illustrate the generation results in Fig. 7. Since only channels that are preserved on both masks are updated, taking the intersection can effectively filter out action-agnostic features, leading to better customization of the actions. In contrast, taking the union may dilute the most representative action features due to the preserved context information.